OpenStudy (anonymous):

Solve, using the substitution method. y – 2x = 8 16 + 4x = 2y

hartnn (hartnn):

y-2x = 8 add 2x on both sides, what u get ?

OpenStudy (anonymous):

y=9+2y

hartnn (hartnn):

now plug in y= 8+2x in 2nd equation

OpenStudy (anonymous):

I thought of doing -2x=8-y

hartnn (hartnn):

that can work, but things will get complicated soon.

OpenStudy (anonymous):

16+4x=2(8+2x)

OpenStudy (anonymous):

I see... I will use the way u told me than

hartnn (hartnn):

yes, simplify 2(8+2x)

OpenStudy (anonymous):

16+4x=16+4x

OpenStudy (anonymous):

O.O IMPOSSIBLE

hartnn (hartnn):

correct. now think logically, when is that true

hartnn (hartnn):

not impossible its like 0=0

OpenStudy (anonymous):

S0 infinite number of solutions

hartnn (hartnn):

you take any value of x or y 16+4x will always be = 16+4x so, yes, infinite solutions :)

OpenStudy (anonymous):

yay! Thanks so much

OpenStudy (mathmale):

Note how one of the given equations is a multiple of the other given equation. That's all the info you need to decide "infinitely many solutions."

OpenStudy (anonymous):

I see, Thanks :)

OpenStudy (mathmale):

:)